OpenStudy (whiskyryan):

Can someone please walk me through this problem:

OpenStudy (whiskyryan):

\[\frac{ \frac{ 1 }{ (x+h)^{2}}-\frac{ 1 }{x ^{2} } }{ h} \]

OpenStudy (mathmale):

Glad to help, but would you please include the instructions for this problem first? This problem is very similar to one that you did a few minutes ago; only the sign of the final result is different. You might want to review your previous question.

OpenStudy (whiskyryan):

It is the same problem... I confused myself and got lost and need to start fresh.... instructions are: simplify the fractional expression.

OpenStudy (mathmale):

1. Identify the LCD. What is it?

OpenStudy (whiskyryan):

(x+h)^2 * x^2?

zepdrix (zepdrix):

yes

zepdrix (zepdrix):

You have a couple of options, 1. You can create a common denominator between the fractions. 2. You can multiply through by the LCM of the denominators. I prefer the second route,\[\large\rm \frac{\frac{1}{(x+h)^2}-\frac{1}{x^2}}{h}\color{royalblue}{\cdot\left(\frac{(x+h)^2\cdot x^2}{(x+h)^2\cdot x^2}\right)}\] if it's too confusing though, you can do the first.

OpenStudy (whiskyryan):

Ok 1 sec... let me work that out...

OpenStudy (whiskyryan):

\[\frac{ x ^{2-(x+h)^{2}} }{ h(x+h)^{2}x ^{2}}\]

zepdrix (zepdrix):

\[\large\rm \frac{x^2-(x+h)^2}{hx^2(x+h)^2}\]Yup, looks good so far. Now expand out the square in the `numerator`.

OpenStudy (whiskyryan):

x^2-x^2+xh+hx+h^2

OpenStudy (whiskyryan):

x^2+x^2+2hx+h^2

zepdrix (zepdrix):

Ok good. But realize that the subtraction in the numerator is being applied to the `entire expansion`.\[\large\rm x^2-(x^2+xh+hx+h^2)\]

OpenStudy (whiskyryan):

oh shoot ... ok i see what i did

zepdrix (zepdrix):

\[\large\rm \frac{x^2-(x^2+2xh+h^2)}{hx^2(x+h)^2}\]This is where we are at so far.

OpenStudy (whiskyryan):

ok... do I exand the (x+h)^2 in the denominator now?

zepdrix (zepdrix):

Nahhh I don't think we want to do that. The whole point of this exercise is to try and cancel out the h in the denominator. Expanding the denominator is just going to make a big mess.

OpenStudy (mathmale):

No. Instead, simplify the numerator. x^2 - x^2 = ?

zepdrix (zepdrix):

Distribute the negative in the numerator, and then try to combine like-terms.

OpenStudy (mathmale):

After you've done that correctly, you can then cancel out the h from numerator and denominator.

OpenStudy (whiskyryan):

so: \[\frac{x^2(-x^2-2xh-h^2)}{hx^2(x+h)^2} \] ?

zepdrix (zepdrix):

After you distribute the negative, drop the brackets,\[\large\rm \frac{x^2-x^2+2xh+h^2}{hx^2(x+h)^2}\]It's important to realize that the operation in front of the brackets shouldn't change from `subtraction` to `multiplication`. That would be weird. That's what it looks like in your expression.

zepdrix (zepdrix):

Woahh I didn't distribute the negative, my bad.

zepdrix (zepdrix):

\[\large\rm \frac{x^2-x^2-2xh-h^2}{hx^2(x+h)^2}\]

OpenStudy (whiskyryan):

ok i see

OpenStudy (mathmale):

Again: Please simplify the numerator. Write out your result. Look for an additional way to reduce the fraction.

OpenStudy (whiskyryan):

\[\frac{-2xh-h^2}{hx^2(x+h)^2}\]

OpenStudy (whiskyryan):

is that right?

zepdrix (zepdrix):

Sorry ran off for a moment :) Looks good so far! You can factor something out of each term in the numerator.

OpenStudy (whiskyryan):

h(-2x+h)

zepdrix (zepdrix):

h(-2x-h)

OpenStudy (whiskyryan):

woops that was a typo... wrote it right on paper

zepdrix (zepdrix):

\[\large\rm\frac{h(-2x-h)}{hx^2(x+h)^2}\]From that point, you should see a cancellation available.

OpenStudy (whiskyryan):

\[\frac{(-2x-h)}{x^2(x+h)^2} \]

zepdrix (zepdrix):

Yayyyyyy good job \c:/ That's pretty much all that needs to be done. What type of math are you doing? In Calculus, we take this difference quotient, and we let this value h approach 0, so it's important that it not show up in the denominator as it did before, that screws things up for us.

OpenStudy (whiskyryan):

Thank you so much for your patience! I am doing pre-calculus... it has been a very long time since I have done math so I am needing to re-learn the basics as I go...

zepdrix (zepdrix):

Ahhh I know that game :) I had to do that a few years back. Math was always my favorite but I had to start back at college Algebra since I hadn't done math in 10 years. It's actually good retaking some of those simple foundational courses. It really really refreshes your mind.